Hello,
I need help with the following problems as I get an error message on my calculator!
its:
1. sqrt x-1=3 ( I don't know the square root symbol!)
2. sqrt x^3=8
3. 3 sqrt x^2=4
4. Is sqrt x^2 power an identity (true for all values of x)?
Thanks so much for your help!! :)

Answer by ankor@dixie-net.com(22740)

1. sqrt(x-1) = 3
:
(Sqrt(x-1))^2 = 3^2; square both sides
x - 1 = 9; gets rid of the radical
x = 9 + 1
x = 10
Check: sqrt(10-1) = 3
:
2. sqrt(x^3) = 8
Square both sides again:
x^3 = 64
x = CubeRt(64)
x = 4
Check: Sqrt(4^3) = 8
:
3. 3*sqrt(x^2) = 4
Square both sides:
(3*sqrt(x^2) = 4^2
9*(x^2) = 16
x^2 = 16/9
x = Sqrt(16/9)
x = +/-4/3
Check 3 * sqrt[(4/3)^2] = 4 (on a calc)
:
4. Is sqrt x^2 power an identity (true for all values of x)?
I think it is